# Large Language Models Are Unconscious of Unreasonability in Math Problems

### v1

@@ Appendix A Prompts for generating UMP
Assistant Prompt As an AI assistant, please assist in generating unreasonable questions based on a given reasonable one. First, identify the variables involved in the problem. Then, modify one variable to render the problem mathematically incorrect. For example, you can make the problem {} Finally, explain how the modification makes the question unreasonable. Remember, the known variables should remain reasonable, and the question should not reveal its unreasonableness solely from the answer. the option in the {}: Undefined Variables: does not provide enough information or variables to solve it. Illogical Scenarios: presents an impossible or illogical situation. Incorrect Assumptions: assumes something that is mathematically incorrect. (For instance, division by zero or fractional individuals in a group, cutting a cake into negative pieces.) Misinterpretation of Units: does not clarify or incorrectly uses units of measurement. Inconsistent Conditions: conditions contradict each other.
Appendix B Examples in UMP benchmark
@@
In this section we will show one unreasonable problem for each category in our benchmark in table 3. ’Answer’ refers to the solution to the original problem. ’New Question’ denotes the artificially generated question that is designed to be unreasonable. Accompanying ’Explanation’ provides rationale for the question’s lack of reason, serving as a guide for evaluators to scrutinize the original problem’s construction. Our intention is that this explanation aids test creators in reviewing and refining their problems to ensure logical consistency.
Appendix C Query format for Pre-validation experiment
@@
Our prompt is as follows: Please decide whether the following statement is reasonable:{d}.Where "d" is the description shown in the following table
@@
Appendix D Example output by different LLM
In this section, we will demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought (CoT) approach compared to our Comparative Cognition Calibration (CCC) method. Tables 4 and LABEL:tab:output_of_qwen-max display the outputs from Chatglm (version 3-6B) and Qwen-Max using two different prompts. It is evident that the outputs generated using our CCC method exhibit marked enhancements

### v2

@@ B.1 Prompt for Semi-automatic framework
Assistant Prompt As an AI assistant, please assist in generating unreasonable questions based on a given reasonable one. First, identify the variables involved in the problem. Then, choose one rule below to modify the problem mathematically incorrect. Finally, explain how the modification makes the question unreasonable. Remember, the known variables should remain reasonable, and the question should not reveal its unreasonableness solely from the answer. Please first state which rule you have chosen and then give the unreasonable question. the rules are here: Undefined Variables: To make the problem does not provide enough information or variables to solve it. Illogical Scenarios: To make the problem presents an impossible or illogical situation. Incorrect Assumptions: To make the problem mathematically incorrect. (For instance, division by zero or fractional individuals in a group, cutting a cake into negative pieces.) Misinterpretation of Units: To make the problem does not clarify or incorrectly uses units of measurement. Inconsistent Conditions: To make the conditions contradict each other. Q:{}
B.2 Why do we need Manual Review
To better test the model’s detection capabilities, we aim to make the unreasonable problems in the benchmark less conspicuous. Initial model-generated questions, particularly within the “Incorrect Assumptions” and “Undefined Variables” categories, were often too evident in their unreasonableness. For instance, some questions will expose the incorrectness directly on the stem, such as the direct appearance of “-3 people” and “x apples” such problematic statements. Therefore, we conducted thorough manual reviews and edited the content. The alterations included changing variables and rephrasing, where necessary, to render the questions superficially reasonable while preserving their underlying unreasonable part. That is why our process requires some kind of manual review.
Appendix C Examples in UMP benchmark
@@
In this section we will show one unreasonable problem for each category in our benchmark in table 3. “Answer” refers to the solution to the original problem. “New Question” denotes the artificially generated question that is designed to be unreasonable. Among them, the part that becomes unreasonable by modification is already represented in bold. Accompanying “Explanation” provides rationale for the question’s lack of reason, serving as a guide for evaluators to scrutinize the original problem’s construction. Our intention is that this explanation aids test creators in reviewing and refining their problems to ensure logical consistency.
Appendix D Query format for Pre-validation experiment
@@
Our prompt is as follows: Please decide whether the following statement is reasonable:{d}.Where “d” is the description shown in the following table
@@
Appendix E Template for CCC
Our CCC template for solving math problems. Initially, we prompt the model with the instruction, “Please solve these problems with a critical thinking.” This is immediately followed by, “Your response should include a step-by-step calculation process.” Subsequently, the model is tasked with assessing the reasonableness of the problem, leading to a conclusion based on this evaluation. This critical juncture is articulated through the guideline, “If the problem is reasonable, please provide the correct answer. If the problem are unreasonable, highlight these issues clearly in your response and provide a succinct explanation.” Finally, the model is instructed to “provide a succinct explanation,” aiming to foster a deeper understanding of the problem-solving process and its outcomes.
CCC template for solving math problems As an AI assistant, you’re tasked with providing accurate solutions to mathematical problems. Please solve these problems with criticism. Your response should include a step-by-step calculation process first and conclude whether the question is reasonable or unreasonable. If the problem is reasonable, please provide the correct answer. If the problem are unreasonable, highlight these issues clearly in your response and provide a succinct explanation.
Appendix F Example output by different LLM
In this section, we will demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought (CoT) approach compared to our CCC method. Tables 4 and 5 display the outputs from Chatglm3-6b and Qwen-Max using two different prompts. It is evident that the outputs generated using our CCC method exhibit marked enhancements.
Appendix G Related Work
G.1 LLMs as Math Problem Solver
G.2 Improve models’ inference ability
Methodologically, a variety of strategies have been developed to augment the model’s reasoning capabilities. For instance, the “chain of thought” approach, introduced by Wei et al. (2022), enables models to generate more accurate solutions by sequentially processing information and emulating human cognitive processes. Complex COT(Fu et al., 2023) prompts models with the most intricate problems and solutions, thereby encouraging the generation of sophisticated responses. Plan-and-Solve (Wang et al., 2023) method enhance model performance by first generating a plan or sequence of steps, which is then executed to arrive at a solution. Yu et al.’s (2023a) output-supervised verifiers refine the model’s capabilities by scrutinizing the generated output and emphasizing the most critical steps within the problem-solving process.